# Fourier neural operator for large eddy simulation of compressible Rayleigh-Taylor turbulence

### v1

@@ I INTRODUCTION
The RANS can provide mean-field information by modeling the Reynolds stresses, which have been widely studied and used in the RT, Richtmyer-Meshkov (RM), and Kelvin-Helmholtz (KH) turbulence. Zhou (2017b); Zhang et al. (2020); Schilling (2021); Morgan (2022) The LES simulates the large-scale motions by modeling the sub-grid scale (SGS) terms, which is more accurate than the RANS and computationally cheaper than DNS. Implicit large eddy simulations (ILES) for solving Euler or Navier-Stokes (NS) equations have been widely used in turbulence mixing,Dimonte et al. (2004) including TURMOIL3D code written by Youngs Youngs (1991, 1994), RTI-3D code developed by Andrews,Andrews (1995) and Miranda code developed at Lawrence Livermore National Laboratory.Cook (2007, 2009); Cook, Cabot, and Miller (2004) Some explicit LES methods based on SGS models have also been used for the study of RT turbulence.Mellado, Sarkar, and Zhou (2005); Nicoud and Ducros (1999); Burton (2011); Mattner, Pullin, and Dimotakis (2004); Zhou et al. (2021); Luo et al. (2023)
@@
With the development of deep learning (ML) technology, neural networks (NNs) have been widely applied to improve or replace traditional computational fluid dynamics (CFD) methods in turbulence simulations.Brunton, Noack, and Koumoutsakos (2020); Duraisamy, Iaccarino, and Xiao (2019)Various methods focus on using NNs to learn closure models for Reynolds stresses in RANS and SGS stresses in LES.Sarghini, De Felice, and Santini (2003); Yuan et al. (2021); Gamahara and Hattori (2017); Guan et al. (2022); Yang et al. (2019); Park and Choi (2021); Ling, Kurzawski, and Templeton (2016); Tabe Jamaat and Hattori (2023); Zhou et al. (2019); Wang et al. (2018a); Xie et al. (2019); Xie, Wang, and E (2020); Xie, Yuan, and Wang (2020); Xu et al. (2023); Beck, Flad, and Munz (2019); Maulik et al. (2019)Deep neural networks (DNNs) exhibit excellent performance in approximating highly nonlinear functions.LeCun, Bengio, and Hinton (2015) Some pure data-driven methods aim to approximate the complete NS equations using DNNs.Lusch, Kutz, and Brunton (2018); Sirignano and Spiliopoulos (2018); Goswami et al. (2022); Mohan et al. (2020); Li, Shu, and Barati Farimani (2024)Once trained, “black-box” NN models can be rapidly deployed on modern computers, achieving much higher efficiency compared to traditional CFD methods. Some studies explored incorporating physical information into DNNs. Cai et al. (2021); Lanthaler, Mishra, and Karniadakis (2022); Karniadakis et al. (2021); Raissi, Perdikaris, and Karniadakis (2019); Wang et al. (2020); Jin et al. (2021); Xu, Zhang, and Wang (2021)
@@
Most DNN architectures aim to learn nonlinear mapping between finite-dimensional Euclidean spacesRaissi, Perdikaris, and Karniadakis (2019); Wu and Xiu (2020); Xu, Zhang, and Zeng (2021). They excel at learning flow fields for specific flow states, but lack sufficient generalization ability for flow parameters, initial conditions, or boundary conditions. Li et al. (2020a) proposed the Fourier neural operator (FNO) method, which can effectively learn the mapping between infinite-dimensional spaces of input-output pairs. The FNO method learns the entire family of PDEs. The FNO method mimics the pseudospectral method, parametrizing the integral kernels in Fourier space, thus directly learning the mapping from function parameters to solutions. Benefiting from its rich expressive power and efficient architecture, the FNO model outperforms previous U-NetChen, Viquerat, and Hachem (2019), TF-NetWang et al. (2020), and ResNetHe et al. (2016). In low Reynolds number two-dimensional (2D) turbulence, the FNO method achieves a prediction error rate of 1%.Li et al. (2020a)  Meng et al. (2023) applied the FNO method to rapid prediction of the dynamic stall of 2D airfoils, also yielding excellent results. Li et al. (2022a) applied the FNO model to LES of three-dimensional (3D) homogeneous isotropic turbulence (HIT), and its predictive performance was significantly better than traditional DSM and DMM models.
@@
After the FNO method was proposed, it was extensively studied, and subsequently, some improved models were developed.Wang et al. (2024); Li et al. (2023a); Deng et al. (2023); Lehmann et al. (2023); Lyu et al. (2023); Peng et al. (2024) In 2D turbulence, Peng, Yuan, and Wang (2022) introduced an attention-enhanced FNO model that can accurately reconstruct the instantaneous structures and statistics of 2D turbulence at high Reynolds numbers, further enhancing the prediction accuracy of the FNO model. Subsequently, they successfully applied it to 3D turbulence by linear attention approximation. Peng et al. (2023) The linear attention coupled FNO method (LAFNO) demonstrates excellent predictive performance in 3D HIT and free shear turbulence. Wen et al. (2022) developed a U-Net enhanced FNO method (U-FNO), which exhibits superior accuracy and efficiency in multi-phase flow problems. You et al. (2022) proposed an implicit FNO method (IFNO), modeling the increment between layers as integral operators to capture long-term dependencies in feature spaces. Combining the advantages of U-FNO and IFNO, Li et al. (2023b) proposed an implicit U-Net enhanced FNO method (IU-FNO) for LES of turbulence. IU-FNO achieves efficient and accurate long-term predictions of turbulence, demonstrating a good performance in 3D HIT, free shear turbulenceLi et al. (2023b), and turbulent channel flowsWang et al. (2024). Guibas et al. (2021) developed an adaptive FNO method (AFNO). AFNO splits high-resolution flow fields into several small local flow fields and uses FNO for block-wise prediction, addressing issues such as low training efficiency and insufficient GPU memory caused by high degrees of freedom. AFNO was employed by NVIDIA in global weather forecasting systems, achieving nearly a five-fold improvement in prediction efficiency compared to traditional methods while maintaining the same prediction accuracy. Li et al. (2024) introduced a transformer-based neural operator (TNO) to achieve precise and efficient predictions for LES of 3D HIT and free shear turbulence. The TNO model can achieve more stable long-term predictions and has fewer parameters than the FNO model.
@@ III The Fourier neural operator
Figure 1: The architecture of Fourier neural operator.
The nonlinear mapping between two infinite-dimensional spaces is denoted as $G^{\dagger}:\mathcal{A}\rightarrow\mathcal{U}$, where $\mathcal{A}=\mathcal{A}\left(D;\mathbb{R}^{d_{a}}\right)$ and $\mathcal{U}=\mathcal{U}\left(D;\mathbb{R}^{d_{u}}\right)$ respectively represent separable Banach spaces valued in $\mathbb{R}^{d_{a}}$ and $\mathbb{R}^{d_{u}}$, with $D\subset\mathbb{R}^{d}$ denoting a bounded open set, and $\mathbb{R}$ representing the real number space.  By constructing a mapping $G:\mathcal{A}\times\Theta\rightarrow\mathcal{U}$ parameterized by $\theta\in\Theta$, the FNO learns an approximation of $G^{\dagger}$. Optimal parameters $\theta^{\dagger}\in\Theta$ can be determined through data-driven methods, so that $G(\cdot,\theta^{\dagger})\approx G^{\dagger}$.Vapnik (1999)
The Fourier neural operator is formulated as an iterative architecture $v_{0}\mapsto v_{1}\mapsto\ldots\mapsto v_{N_{f}}$, where $v_{j}$ for $j=0,1,\ldots,N_{f}$ is a sequence of functions each taking values in $\mathbb{R}^{d_{v}}$. Li et al. (2020b) The architecture of FNO is illustrated in Fig. 1, comprising three main steps:Li et al. (2020a, 2022a)
(1) The input $a\in\mathcal{A}$ is lifted to a higher dimensional representation by a local transformation $P$, denoted as $v_{0}(x)=P(a(x))$, parameterized by a shallow fully connected neural network.
(2) The high-dimensional $v_{0}(x)$ is updated through $N_{f}$ iterations, each iteration represented as
@@
where $\mathcal{K}:\mathcal{A}\times\Theta_{\mathcal{K}}\rightarrow\mathcal{L}\left(\mathcal{U}\left(D;\mathbb{R}^{d_{v}}\right),\mathcal{U}\left(D;\mathbb{R}^{d_{v}}\right)\right)$ maps to bounded linear operators on $\mathcal{U}\left(D;\mathbb{R}^{d_{v}}\right)$ and is parameterized by $\phi\in\Theta_{\mathcal{K}}$, $W:\mathbb{R}^{d_{v}}\rightarrow\mathbb{R}^{d_{v}}$ is a linear transformation, and $\sigma:\mathbb{R}\rightarrow\mathbb{R}$ is a nonlinear activation function whose action is defined component-wise.
(3) The output $u\in\mathcal{U}$ is the projection of $v_{N_{f}}(x)$ by the local transformation $Q:\mathbb{R}^{d_{v}}\rightarrow\mathbb{R}^{d_{u}}$, denoted as $u(x)=Q\left(v_{N_{f}}(x)\right)$, parameterized by a fully connected layer.Li et al. (2020a)
@@
$\left(\mathcal{K}(\phi)v_{t}\right)(x)=\mathcal{F}^{-1}\left(R_{\phi}\cdot\left(\mathcal{F}v_{t}\right)\right)(x),\quad\forall x\in D,$ (24)
where $R_{\phi}$ is the Fourier transform of a periodic function $\mathcal{K}:\bar{D}\rightarrow\mathbb{R}^{d_{v}\times d_{v}}$, parameterized by $\phi\in\Theta_{\mathcal{K}}$. The finite-dimensional parameterization by truncating the Fourier series at a maximal number of modes $k_{\max}=\left|Z_{k_{\max}}\right|=\mid\left\{k\in\mathbb{Z}^{d}:\left|k_{j}\right|\leq k_{\max,j}\right.,\left.j=1,\ldots,d\right\}\mid$. $R_{\phi}$ is parameterized as a complex tensor $({k_{\max}\times d_{v}\times d_{v}})$, which contains a collection of truncated Fourier modes $R_{\phi}\in\mathbb{C}^{k_{\max}\times d_{v}\times d_{v}}$. The domain $D$ is assumed to be discretized with $n\in\mathbb{N}$ points, where $v_{t}\in\mathbb{R}^{n\times d_{v}}$.Li et al. (2020a)  $\mathcal{F}\left(v_{t}\right)\in\mathbb{C}^{k_{\max}\times d_{v}}$ can be obtained by truncating the higher modes. Therefore, $\left(R_{\phi}\cdot\left(\mathcal{F}v_{t}\right)\right)_{k,l}=\sum_{j=1}^{d_{v}}R_{\phi k,l,j}\left(\mathcal{F}v_{t}\right)_{k,j},\ k=1,\ldots,k_{\max}$, and $\ j=1,\ldots,d_{v}$.
@@ IV Dataset description and FNO model settings
When using neural network methods for flow field prediction, we need to train the model with a large amount of data. In this study, the data comes from numerical simulations of compressible RT turbulence in a $1^{2}\times 2$ rectangular box with a grid scale of $\Delta_{x}=1/128$. Thus, there are a total of $128^{2}\times 256$ grid points. The Reynolds number is $Re=10000$. The data in current numerical simulations with $128^{2}\times 256$ grids is referred to as ground truth (GT) data. The numerical simulations with $128^{2}\times 256$ grid points are not sufficient to reach the spectral resolution of the Nyquist limit at $Re=10000$Lele (1992). For the lower order statistics considered in this study, the results of ground truth are close to the DNS results. When training the model, we need to coarsen the high-fidelity data, selecting uniformly $32^{2}\times 64$ grids of the flow field data. We refer to the flow field with the coarse grids as the filtered ground truth (fGT) flow field. The Atwood number and stratification parameter are $A_{t}=0.5$ and $Sr=1.0$, respectively. Since the central plane has only $32^{2}$ grids, the wave number range of the initial disturbance is $5\leq k\leq 10$, ensuring sufficient resolution for the initial disturbance.
@@
In each case, the total numerical simulation time is $3\tau$, where the dimensionless time $\tau=\sqrt{{L_{r}}/(A_{t}g)}$, and the time step is $\Delta t=0.001\tau$ in numerical simulations with $128^{2}\times 256$ grids. Here, we only use data from the first two dimensionless time during the training of the FNO model, resulting in a total of 2001 time points for the filtered flow field data. The time step used by the FNO model is much larger than that of numerical simulations with $128^{2}\times 256$ grids, which is called the time step of the FNO model, $\Delta t_{FNO}$. We test $\Delta t_{FNO}=20\Delta t$, $40\Delta t$, $60\Delta t$, $80\Delta t$. The results show that the prediction performance of the FNO model is the best for $\Delta t_{FNO}=60\Delta t$ in the a posteriori tests. If the time step of the FNO model is too large, it might miss some flow field structures. But if $\Delta t_{FNO}$ is too small, there would be too much high-frequency information in the flow field, leading to divergence of the FNO model in the a posteriori tests. Although the FNO model trained with a smaller $\Delta t_{FNO}$ accurately predicts the flow field in the first few steps, it is difficult to obtain convergent long-term predictive results. Therefore, we actually use only $(300\times 34)$ sets of data during the training of the FNO model. In the a posteriori tests, we need to predict for a longer time, $3\tau$, to examine the generalization performance of the FNO model over a longer time.
@@
In summary, the training data for the FNO model consist of $[N_{c}\times N_{t}\times N_{x}\times N_{y}\times N_{z}\times N_{v}]$ samples, where $N_{c}=300$ represents the number of independent cases, $N_{t}=34$ denotes the number of FNO time steps, $N_{x}=32,N_{y}=32,N_{z}=64$ represent the number of grid points in three coordinate directions, and $N_{v}=6$ indicates the number of physical variables. Therefore, there are a total of $[300\times 34\times 32\times 32\times 64\times 6]$ samples, with 80% of the data used for training and 20% for testing. The dimensionless physical variables of the $n$-th FNO time step are denoted as $\mathbf{f}^{*}_{n}$. In the FNO model, the physical fields of the previous five FNO time steps $(\mathbf{f}^{*}_{1},\mathbf{f}^{*}_{2},\mathbf{f}^{*}_{3},\mathbf{f}^{*}_{4},\mathbf{f}^{*}_{5})$ are taken as input, and the physical variables at the sixth FNO time step $\mathbf{f}^{*}_{6,pre}$ are taken as output, as shown in Fig. 1. Then, the flow fields of the 7-th time step $\mathbf{f}^{*}_{7,pre}$ are predicted using $(\mathbf{f}^{*}_{2},\mathbf{f}^{*}_{3},\mathbf{f}^{*}_{4},\mathbf{f}^{*}_{5},\mathbf{f}^{*}_{6,pre})$, and so on. Therefore, we need to predict 46 steps to obtain data at $3\tau$ in the a posteriori tests. During the training of the FNO model, we use the dimensionless physical fields $\mathbf{f}^{*}$. In the a posteriori tests, we also obtain dimensionless physical fields, and then we need to use the rms values of the previous time step to obtain the true flow field ${f}_{n}={f}^{*}_{n}\times f^{rms}_{n-1}(0)~$ at the current time step.

### v2

@@ I INTRODUCTION
The RANS can provide mean-field information by modeling the Reynolds stresses, which have been widely studied and used in the RT, Richtmyer-Meshkov (RM), and Kelvin-Helmholtz (KH) turbulence. Zhou (2017b); Zhang et al. (2020); Schilling (2021); Morgan (2022); Xiao, Zhang, and Tian (2021); Kokkinakis, Drikakis, and Youngs (2020) The LES simulates the large-scale motions by modeling the sub-grid scale (SGS) terms, which is more accurate than the RANS and computationally cheaper than DNS. Implicit large eddy simulations (ILES) for solving Euler or Navier-Stokes (NS) equations have been widely used in turbulence mixing,Dimonte et al. (2004) including TURMOIL3D code written by Youngs Youngs (1991, 1994), RTI-3D code developed by Andrews,Andrews (1995) and Miranda code developed at Lawrence Livermore National Laboratory.Cook (2007, 2009); Cook, Cabot, and Miller (2004) Some explicit LES methods based on SGS models have also been used for the study of RT turbulence.Mellado, Sarkar, and Zhou (2005); Nicoud and Ducros (1999); Burton (2011); Mattner, Pullin, and Dimotakis (2004); Zhou et al. (2021); Luo et al. (2023)
@@
With the development of deep learning (ML) technology, neural networks (NNs) have been widely applied to improve or replace traditional computational fluid dynamics (CFD) methods in turbulence simulations.Brunton, Noack, and Koumoutsakos (2020); Duraisamy, Iaccarino, and Xiao (2019)Various methods focus on using NNs to learn closure models for Reynolds stresses in RANS and SGS stresses in LES.Sarghini, De Felice, and Santini (2003); Yuan et al. (2021); Gamahara and Hattori (2017); Guan et al. (2022); Yang et al. (2019); Park and Choi (2021); Ling, Kurzawski, and Templeton (2016); Tabe Jamaat and Hattori (2023); Wang et al. (2018a); Xie et al. (2019); Xie, Wang, and E (2020); Xie, Yuan, and Wang (2020); Xu et al. (2023); Beck, Flad, and Munz (2019); Maulik et al. (2019); Huang, Yang, and Kunz (2019); Huang and Yang (2021); Qi et al. (2022); Yin et al. (2022); Meng, Jiang, and Wang (2023); Li et al. (2021)Deep neural networks (DNNs) exhibit excellent performance in approximating highly nonlinear functions.LeCun, Bengio, and Hinton (2015) Some pure data-driven methods aim to approximate the complete NS equations using DNNs.Lusch, Kutz, and Brunton (2018); Sirignano and Spiliopoulos (2018); Goswami et al. (2022); Mohan et al. (2020); Li, Shu, and Barati Farimani (2024)Once trained, “black-box” NN models can be rapidly deployed on modern computers, achieving much higher efficiency compared to traditional CFD methods. Some studies explored incorporating physical information into DNNs. Cai et al. (2021); Lanthaler, Mishra, and Karniadakis (2022); Karniadakis et al. (2021); Raissi, Perdikaris, and Karniadakis (2019); Wang et al. (2020); Jin et al. (2021); Xu, Zhang, and Wang (2021)
@@
Most DNN architectures aim to learn nonlinear mapping between finite-dimensional Euclidean spacesRaissi, Perdikaris, and Karniadakis (2019); Wu and Xiu (2020); Xu, Zhang, and Zeng (2021). They excel at learning flow fields for specific flow states, but lack sufficient generalization ability for flow parameters, initial conditions, or boundary conditions. Li et al. (2020a) proposed the Fourier neural operator (FNO) method, which can effectively learn the mapping between infinite-dimensional spaces of input-output pairs. The FNO method learns the entire family of PDEs. The FNO method mimics the pseudospectral method, parametrizing the integral kernels in Fourier space, thus directly learning the mapping from function parameters to solutions. Benefiting from its rich expressive power and efficient architecture, the FNO model outperforms previous U-NetChen, Viquerat, and Hachem (2019), TF-NetWang et al. (2020), and ResNetHe et al. (2016). In low Reynolds number two-dimensional (2D) turbulence, the FNO method achieves a prediction error rate of 1%.Li et al. (2020a)  Meng, Jiang, and Wang (2023) applied the FNO method to rapid prediction of the dynamic stall of 2D airfoils, also yielding excellent results. Li et al. (2022a) applied the FNO model to LES of three-dimensional (3D) homogeneous isotropic turbulence (HIT), and its predictive performance was significantly better than traditional DSM and DMM models.
@@
After the FNO method was proposed, it was extensively studied, and subsequently, some improved models were developed.Wang et al. (2024); Li et al. (2023a); Deng et al. (2023); Lehmann et al. (2023); Lyu et al. (2023); Peng et al. (2024); Qin et al. (2024) In 2D turbulence, Peng, Yuan, and Wang (2022) introduced an attention-enhanced FNO model that can accurately reconstruct the instantaneous structures and statistics of 2D turbulence at high Reynolds numbers, further enhancing the prediction accuracy of the FNO model. Subsequently, they successfully applied it to 3D turbulence by linear attention approximation. Peng et al. (2023) The linear attention coupled FNO method (LAFNO) demonstrates excellent predictive performance in 3D HIT and free shear turbulence. Wen et al. (2022) developed a U-Net enhanced FNO method (U-FNO), which exhibits superior accuracy and efficiency in multi-phase flow problems. You et al. (2022) proposed an implicit FNO method (IFNO), modeling the increment between layers as integral operators to capture long-term dependencies in feature spaces. Combining the advantages of U-FNO and IFNO, Li et al. (2023b) proposed an implicit U-Net enhanced FNO method (IU-FNO) for LES of turbulence. IU-FNO achieves efficient and accurate long-term predictions of turbulence, demonstrating a good performance in 3D HIT, free shear turbulenceLi et al. (2023b), and turbulent channel flowsWang et al. (2024). Guibas et al. (2021) developed an adaptive FNO method (AFNO). AFNO splits high-resolution flow fields into several small local flow fields and uses FNO for block-wise prediction, addressing issues such as low training efficiency and insufficient GPU memory caused by high degrees of freedom. AFNO was employed by NVIDIA in global weather forecasting systems, achieving nearly a five-fold improvement in prediction efficiency compared to traditional methods while maintaining the same prediction accuracy. Li et al. (2024) introduced a transformer-based neural operator (TNO) to achieve precise and efficient predictions for LES of 3D HIT and free shear turbulence. The TNO model can achieve more stable long-term predictions and has fewer parameters than the FNO model.
@@ III The Fourier neural operator
Figure 1: The architecture of Fourier neural operator for 3D RT turbulence prediction. The middle is the FNO architecture, the lower left corner shows the input of the model, and the lower right corner shows the output. Here, $N_{x}=32,N_{y}=32$ and $N_{z}=64$ respectively represent grid resolutions in three coordinate directions, $N_{v}=6$ indicates the number of physical variables, $N_{I}=5$ is the number of time-nodes of input, and $N_{c}=3$ is the number of dimensions.
The nonlinear mapping between two infinite-dimensional spaces is denoted as $G^{\dagger}:\mathcal{A}\rightarrow\mathcal{U}$, where $\mathcal{A}=\mathcal{A}\left(D;\mathbb{R}^{d_{a}}\right)$ and $\mathcal{U}=\mathcal{U}\left(D;\mathbb{R}^{d_{u}}\right)$ respectively represent separable Banach spaces valued in $\mathbb{R}^{d_{a}}$ and $\mathbb{R}^{d_{u}}$, with $D\subset\mathbb{R}^{d}$ denoting a bounded open set, $\mathbb{R}$ representing the real number space. $\mathbb{R}^{d_{a}}$ and $\mathbb{R}^{d_{u}}$ are the value sets of input $a(x)$ and output $u(x)$, respectively, where $d_{a}$ and $d_{u}$ are the dimensions of $a(x)$ and $u(x)$, respectively. The Fourier neural operators learn an approximation of $G^{\dagger}$ by constructing a mapping $G:\mathcal{A}\times\Theta\rightarrow\mathcal{U}$ parameterized by $\theta\in\Theta$. Optimal parameters $\theta^{\dagger}\in\Theta$ can be determined through data-driven optimization methods, so that $G(\cdot,\theta^{\dagger})\approx G^{\dagger}$.Vapnik (1999)
The Fourier neural operator employs iterative architectures $v_{0}\mapsto v_{1}\mapsto\ldots\mapsto v_{N_{f}}$, where $v_{j}$ for $j=0,1,\ldots,N_{f}$ is a sequence of functions each taking values in $\mathbb{R}^{d_{v}}$. Here, $\mathbb{R}^{d_{v}}$ is the real numbers sets of $v_{t}(x)$, and $d_{v}$ is the dimension of $v_{t}(x)$.Li et al. (2020b) The architecture of FNO is illustrated in Fig. 1, comprising three main steps:Li et al. (2020a, 2022a)
(1) The input $a\in\mathcal{A}$ is lifted to a higher dimensional representation by a local transformation $P$, denoted as $v_{0}(x)=P\left(a(x)\right)$, parameterized by a shallow fully connected neural network.
(2) The high-dimensional $v_{0}(x)$ is updated through $N_{f}$ iterations, and each iteration is represented as
@@
where $\sigma:\mathbb{R}\rightarrow\mathbb{R}$ is a nonlinear activation function whose action is defined component-wise, $W:\mathbb{R}^{d_{v}}\rightarrow\mathbb{R}^{d_{v}}$ is a linear transformation, and $\mathcal{K}:\mathcal{A}\times\Theta_{\mathcal{K}}\rightarrow\mathcal{L}\left(\mathcal{U}\left(D;\mathbb{R}^{d_{v}}\right),\mathcal{U}\left(D;\mathbb{R}^{d_{v}}\right)\right)$ maps to bounded linear operators on $\mathcal{U}\left(D;\mathbb{R}^{d_{v}}\right)$ and is parameterized by $\phi\in\Theta_{\mathcal{K}}$.
(3) The output $u\in\mathcal{U}$ is the projection of $v_{N_{f}}(x)$ by the local transformation $Q:\mathbb{R}^{d_{v}}\rightarrow\mathbb{R}^{d_{u}}$, denoted as $u(x)=Q\left(v_{N_{f}}(x)\right)$, parameterized by a fully connected layer.
@@
$\left(\mathcal{K}(\phi)v_{t}\right)(x)=\mathcal{F}^{-1}\left\{R_{\phi}\cdot\mathcal{F}\left\{v_{t}\right\}\right\}(x),\quad\forall x\in D,$ (24)
where $R_{\phi}$ is the Fourier transform of a periodic function $\mathcal{K}:\bar{D}\rightarrow\mathbb{R}^{d_{v}\times d_{v}}$, parameterized by $\phi\in\Theta_{\mathcal{K}}$. $R_{\phi}$ is parameterized as a complex tensor $({k_{\max}\times d_{v}\times d_{v}})$, which contains a collection of truncated Fourier modes $R_{\phi}\in\mathbb{C}^{k_{\max}\times d_{v}\times d_{v}}$. Here, $\mathbb{C}$ denotes the complex space. The domain $D$ is assumed to be discretized with $n\in\mathbb{N}$ points, where $v_{t}\in\mathbb{R}^{n\times d_{v}}$. The finite-dimensional parameterization by truncating the Fourier series at a maximal number of modes $k_{\max}=\left|Z_{k_{\max}}\right|=\mid\left\{k\in\mathbb{Z}^{d}:\left|k_{j}\right|\leq k_{\max,j}\right.,\left.j=1,\ldots,d\right\}\mid$$\mathcal{F}\left\{v_{t}\right\}\in\mathbb{C}^{k_{\max}\times d_{v}}$ can be obtained by truncating the higher modes. Therefore, $(R_{\phi}\cdot\mathcal{F}\left\{v_{t}\right\})_{k,l}=\sum_{j=1}^{d_{v}}R_{\phi k,l,j}\left(\mathcal{F}\left\{v_{t}\right\}\right)_{k,j}$, $k=1,\ldots,k_{\max}$, and $j=1,\ldots,d_{v}$.Li et al. (2020a, 2022a)
@@ IV Dataset description and FNO model settings
The previous section introduced the architecture of Fourier neural operator. In this section, we will introduce the input–output pairs shown in the lower left and lower right corners in Fig. 1, as well as the parameters used in the FNO model in 3D compressible RT turbulence.
When using neural network methods for flow field prediction, we need to train the model with a large amount of data. In this study, the data comes from numerical simulations of compressible RT turbulence in a $1^{2}\times 2$ rectangular box with a grid scale of $\Delta_{x}=1/128$. Thus, there are a total of $128^{2}\times 256$ grid points. The Reynolds number is $Re=10000$. The data in current numerical simulations with $128^{2}\times 256$ grids is referred to as ground truth (GT) data. The numerical simulations with $128^{2}\times 256$ grid points are not sufficient to reach the spectral resolution of the Nyquist limit at $Re=10000$Lele (1992). For the lower-order statistics considered in this study, the results of ground truth are close to the DNS results. The comparison between $512^{2}\times 1024$, $256^{2}\times 512$ and $128^{2}\times 256$ grids is provided in Appendix A. When training the model, we need to downsample the high-fidelity data into the coarse grid, selecting uniformly $32^{2}\times 64$ grids of the flow field data. We refer to the flow field with the coarse grids as the filtered ground truth (fGT) flow field. The Atwood number and stratification parameter are $A_{t}=0.5$ and $Sr=1.0$, respectively. Since the central plane has only $32^{2}$ grids, the wave number range of the initial disturbance is $5\leq k\leq 10$, ensuring sufficient resolution for the initial disturbance.
@@
In each case, the total numerical simulation time is $3\tau$, where the dimensionless time $\tau=\sqrt{{L_{r}}/(A_{t}g)}$, and the time step is $\Delta t=0.001\tau$ in numerical simulations with $128^{2}\times 256$ grids. Here, we only use data from the first two dimensionless time during the training of the FNO model, resulting in a total of 2001 time points for the filtered flow field data. The FNO model can directly predict the flow field over a large time intervalLi et al. (2020a, 2022a). The time step used by the FNO model, $\Delta t_{\rm{FNO}}$, is much larger than that of numerical simulations with $128^{2}\times 256$ grids. We test $\Delta t_{\rm{FNO}}=20\Delta t$, $40\Delta t$, $60\Delta t$, $80\Delta t$. The results show that the prediction performance of the FNO model is the best for $\Delta t_{\rm{FNO}}=60\Delta t$ ($\Delta t_{\rm{FNO}}=0.06\tau$) in the a posteriori tests. We infer that FNO models with large time intervals can filter small-scale structures with high frequency. If the time step of the FNO model is too large, the prediction might miss some flow field structures. If $\Delta t_{\rm{FNO}}$ is too small, the prediction gives rise to the situation with more high-frequency structures in the flow field. High frequency structures are more difficult to learn and lead to more instability in the a posteriori tests. Although the FNO model trained with a smaller $\Delta t_{\rm{FNO}}$ accurately predicts the flow field in the first few steps, it is difficult to obtain convergent long-term predictive results. Therefore, we actually use only $(300\times 34)$ sets of data during the training of the FNO model. In the a posteriori tests, we need to predict for a longer time, $3\tau$, to examine the generalization performance of the FNO model over a longer time.
@@
In summary, the training data for the FNO model consist of $[N_{c}\times N_{t}\times N_{x}\times N_{y}\times N_{z}\times N_{v}]$ samples, where $N_{c}=300$ represents the number of independent cases, $N_{t}=34$ denotes the number of FNO time steps, $N_{x}=32,N_{y}=32,N_{z}=64$ represent the number of grid points in three coordinate directions, and $N_{v}=6$ indicates the number of physical variables. Therefore, there are a total of $[300\times 34\times 32\times 32\times 64\times 6]$ samples, with 80% of the data used for training and 20% for testing. The dimensionless physical variables of the $n$-th FNO time step are denoted as $\mathbf{f}^{*}_{n}$. In the FNO model, the physical fields of the previous five FNO time steps $(\mathbf{f}^{*}_{n},\mathbf{f}^{*}_{n+1},\mathbf{f}^{*}_{n+2},\mathbf{f}^{*}_{n+3},\mathbf{f}^{*}_{n+4})$ are taken as input, which is similar to the situations of 3D incompressible HIT, free shear turbulenceLi et al. (2022a, 2023b), and turbulent channel flowsWang et al. (2024). If the number of history steps is small, the prediction results of the FNO model would be not so accurate. The physical variables at the sixth FNO time step $\mathbf{f}^{*}_{n+5,pre}$ are taken as output, as shown in Fig. 1. Then, the flow fields of the 7-th time step $\mathbf{f}^{*}_{n+6,pre}$ are predicted using $(\mathbf{f}^{*}_{n+1},\mathbf{f}^{*}_{n+2},\mathbf{f}^{*}_{n+3},\mathbf{f}^{*}_{n+4},\mathbf{f}^{*}_{n+5,pre})$, and so on. Therefore, we need to predict 46 steps to obtain data at $3\tau$ in the a posteriori tests. During the training of the FNO model, we use the dimensionless physical fields $\mathbf{f}^{*}$. In the a posteriori tests, we also obtain dimensionless physical fields, and then we need to use the rms values of the previous time step to obtain the true flow field ${f}_{n}={f}^{*}_{n}\times f^{rms}_{n-1}(0)~$ at the current time step.